OpenStudy (gschibby):

Limit when x->2 (x-2)/(sqrt(x^2-3)-(x/2))

OpenStudy (anonymous):

use L'hopitals rule

OpenStudy (anonymous):

differentiate numerator: (1/2)(1/Sqrt[x+2])

OpenStudy (gschibby):

That's the problem, the task is "solve WITHOUT usin l'Hopitals" :/

OpenStudy (anonymous):

differentiate denominator: 1

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

limit becomes: (1/2)(1/Sqrt[7+2]) = 1/2(1/3) = 1/6

OpenStudy (gschibby):

I think I'm supposed to do it something like this:

OpenStudy (gschibby):

where you multiply the denominator and nominator so as to get a (x-2) both above and below the line. I just can't seem to get there :P

OpenStudy (anonymous):

@***[ISURU]***